OpenStudy (anonymous):

In the xy-plane, the line x+y=k, where k is a constant, is tangent to the graph of y=x^2+3x+1 what is the value of k?

OpenStudy (anonymous):

How would one approach this problem

OpenStudy (anonymous):

The derivative of a function represents the tangent line of its curve. This is the key to this problem.

OpenStudy (anonymous):

How would someone apply that concept to the first equation given?

OpenStudy (anonymous):

Or rather how would one set up the problem?

OpenStudy (anonymous):

Well... The question said that the line y=k-x is the tangent of y=x^2+3x+1.. What do you think that means considering the concept I wrote above?

OpenStudy (anonymous):

Y prime of x^2+3x+1 is equal to y= k-x ...?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Oh !!! I know what I need to do. Thank you so much!

OpenStudy (anonymous):

I am sure you do!!